# Quasiperiodicity protects quantized transport in disordered systems without gaps

### v1

@@ Appendix B Critical disorder strengths for first-order bands
We show here how to use the configuration space pictures to derive the critical disorder strength for the the $3$ bands defined by the $n=\pm 1$ gaps, in the case $\beta>1/2$. Note that the general idea can be applied to any band, and any value of $\beta>1/2$. The cases $\beta<1/2$ can be obtained by sending $\beta\rightarrow 1-\beta$ and inverting the sign of all Chern numbers.
@@
Let us consider $3$ nearest neighbors, which we label with indices $i=-1,0,1$, each of them being offset by an onsite disorder of strength $w_{i}$. The resonance condition for the $-1$ and 00 sites is given by
@@
while the resonant condition between sites 00 and $1$ is given by
$2V\cos{(\varphi_{0,1})}+w_{0}=2V\cos{(\varphi_{0,1}+2\pi\beta)}+w_{1}$ (21)
@@
As discussed in Section III, the breakdown of quantized transport occurs once $\Delta\varphi=\varphi_{-1,0}-\varphi_{0,1}=0$. For bounded local disorder this becomes possible as soon as the disorder is strong enough to shift the individual resonances to $\varphi_{-1,0}=\varphi_{0,1}=0$. This condition leads to the following critical disorder strength for the top and bottom $C=+1$ bands:
@@
The same reasoning can be applied to the middle $C=-2$ band, with the only difference that breakdown happens when the resonant angles meet halfway at $\varphi_{left}=\varphi_{right}=\pi/2$, leading to a different critical disorder strength
$W_{crit}^{C=-2}$ $=2V|\sin{\left(\pi\beta\right)}\cos{\left(\pi\beta\right)}|$
$=2V|\sin{\left(2\pi\beta\right)}|/2$ (25)
We stress that these expressions do not depend on the tunneling strength $J$, and hence do not depend on the size of the gaps. This explains the stability of the quantized current for disorder strengths that are larger than the size of the clean gaps.
In addition, we note that the same reasoning can be used to derive the critical disorder strength for different kinds of bounded spatial disorder. For instance, in the presence of random phase disorder in onsite energies of the form $\epsilon_{i}=2V\cos{(2\pi\beta i+\varphi+W_{i})}$ ), the above argument can be directly transposed to conclude that the critical disorder strength required to break the quantized pumping of a given band is directly proportional to the angle (and hence the number of states) in that band:
$W_{crit}^{C=+1}=\pi|1-\beta|$ (26)
$W_{crit}^{C=-2}=\pi|\beta-\frac{1}{2}|$ (27)
In both cases, the topological protection increases with the number of states in the considered energy bands of the clean system and does not rely on a spectral gap condition, which is in stark contrast with the usual intuition of gap-protected topological properties.
Appendix C Critical disorder strengths for second-order central band
The perturbation analysis from Eq. 24 can be extend to second order by looking for the hybridisation of pairs of resonant next-nearest neighbors, mediated by an effective tunneling amplitude (given by $J_{eff}^{(2)}\propto J^{2}/V$). The corresponding resonance condition is given by:
[…]
which causes the opening of two gaps of size $\Delta E^{(2)}/\propto J^{2}/V$, at energies $E^{(\pm 2)}=\pm 2V\cos{4\pi\beta}$ ). These resonances, combined with the first order resonances, split the spectrum in $5$ second-order bands (Fig. 1 c). If the pumping rate is sufficiently slow to allow adiabatic hoppings between next-nearest neighbors, the Thouless pump becomes sensitive to these smaller gaps.
The critical disorder strength required to perturb the quantized current in the $C=+4$ band can be deduced using the argument shown in Fig. 4, with the angles corresponding to the $C=+4$ band. In this case, the angle spanned by the band in configuration space is twice $\Delta\varphi^{C=+4}=2\pi(\frac{3}{4}-\beta)$. Let us consider $3$ next nearest-neighbors, labelled $-1,0,1$. The disorder can shift these resonance by (at most) an angle:
$\varphi_{-1,0}=2\pi(1-\beta)+\arcsin{\left(\frac{w_{0}-w_{-1}}{4V\sin{\left(2\pi\beta\right)}}\right)}$ ) (29)
$\varphi_{0,1}=-2\pi(1-\beta)+\arcsin{\left(\frac{w_{1}-w_{0}}{4V\sin{\left(2\pi\beta\right)}}\right)}$ ) (30)
which leads to a breakdown of the quantized current when the resonances meet at $\varphi=\pi/2$, at the critical disorder strength
$W_{crit}^{C=+4}=2V|\sin{(2\pi\beta)}\cos{(2\pi\beta))}|\,,$ (31)
which shows excellent agreement with numerical simulations (Fig. 3 j).

### v3

@@ Appendix B Estimation of critical disorder strengths for first-order $C=+1$ bands
We use the configuration space pictures to estimate the critical disorder strength for the the $2$ first-order bands with $C=+1$, defined by the $n=\pm 1$ gaps, in the case $\beta>1/2$. Note, the cases $\beta<1/2$ can be obtained by sending $\beta\rightarrow 1-\beta$ and inverting the sign of all Chern numbers.
@@
Let us consider $3$ consecutive sites with indices $i=-1,0,1$, each being offset by an onsite disorder of strength $w_{i}$. The phase $\varphi_{-1,0}$, where the $-1$ and 00 sites are resonant, is given by
@@
and the corresponding condition for sites 00 and $1$ is
$2V\cos{(\varphi_{0,1})}+w_{0}=2V\cos{(\varphi_{0,1}+2\pi\beta)}+w_{1}.$ (22)
@@
for the top $C=+1$ band and
$\varphi_{-1,0}=\pi\beta+\arcsin{\left(\frac{w_{0}-w_{-1}}{4V\sin{\left(\pi\beta\right)}}\right)}$ ) (25)
$\varphi_{0,1}=-\pi\beta+\arcsin{\left(\frac{w_{1}-w_{0}}{4V\sin{\left(\pi\beta\right)}}\right)}$ ) (26)
for the bottom $C=+1$ band.
As discussed in Section III, the breakdown of quantized transport occurs once $\Delta\varphi=\varphi_{-1,0}-\varphi_{0,1}=0$. For bounded local disorder this becomes possible as soon as the disorder is strong enough to shift the individual resonances to $\varphi_{-1,0}=\varphi_{0,1}=0$. This condition leads to the following estimate for the critical disorder strength for the top and bottom $C=+1$ bands: